A piece of 14-gauge copper wire 0. 163cm in diameter is 1 4.0m long. (a) What is its resistance? (b) If a potential


difference (V) of 1 .00 V is applied across the wire, what current flows in it? (c) What is the electric field in the


wire?


1
Expert's answer
2021-05-10T08:27:32-0400

(a) Find the resistance of copper with resistivity 1.724·10-8 Ohm·m:


"R=\frac{\rho L}{A}=\frac{4\rho L}{\pi d^2}=0.116\space\Omega."

(b) The current is


"I=\frac VR=8.65\text{ A}."

(c) The electric field inside the wire is zero.